KTeV Shift Summary - DAY 11/6/99 (Sat.)

Earl C. Swallow (earls@elmhurst.edu)
Sat, 6 Nov 1999 15:43:04 -0600 (CST)

Shift Summary for DAY shift, 11/06/99 (Sat.)

On Shift: Hogan Nguyan and Earl Swallow

Hours of beam available at all intensities: 3.8 (out of 8.5)
Hours of beam available at nominal intensity: 3.7
Hours of E799 data (special runs not included): 3.7
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:

Electrical short in extraction septum and lots of left bend trips.

Runs ended this shift:

14914 -- BEAM_799
14915 -- nothing
14916 -- LaserScan ("excellent")

Tapes brought to FCC this shift:

UPA,B,C100-102, UPS034 -- Runs 14906-14911
UPA,B,C103-105, UPS035 -- Runs 14913-14916

Accesses during shift:


Overview of shift:

Smooth data taking (with good time structure) WHEN we had beam.
Unfortunately beam was off for more than half of shift.

Description of Shift:

14914 07:00 DAY Shift Begins -- smooth running in progress.

DAY Shift tasks (from BobT): (1) Compare "yields"
(livetime*SEM) for 2-turn and 3-turn injection.
(2) Change HV for 6 accidental counter channels
from 1295 V to 1250 V at end of this run.

07:54 Online Split requests tape mount: UPF027 -->
KTeV2/tps2d2. Problem: KTeV2/tps2d2 is in use for
regular DAQ output. Contact Sydney who will fit it
-- probably just a residual configuration from the
shut-sown period.
602 08:00 Spills 602-603 corrupted by accidental setting of
accidental counter JHV to zero (0).

Hogan notes that DC current is MUCH smoother with
2-turn injection.

~617 08:28 Beam OFF -- short in one of our extraction septa.
Experts coming in, etc.

08:54 End Run 14914 -- BEAM_799.

09:05 Accidental counter HVs --> 1250 V.

14916 09:35 Begin Run 14916 -- LaserScan (there is no Run 14915).

Online Split requests tape mount: UPF027 -->
KTEV4/tps3d2 -- Sydney has fixed it.

10:17 End Run 14916 -- LaserScan (Hogan says "excellent").

14917 11:15 Begin Run 14917 -- BEAM_799
(but no beam from Tevatron).

12:05 Hogan adjusts Laser filter wheel to increase light

12:28 Beam returns -- will start at low intensity and work
their way up. It soon looks very good.

12:47 For 2-turn operation, SEM = 8.8E12, Livetime = 0.648.
At the moment, MC is doing tests, so we get almost
all of beam.

Hogan looks at online histograms:
1. L1 timing for RC10 is ~6 ns different than
reference plots
2. Timing for V0,V1 >= 2 is shifted by ~1 ns!
(Sydney already notified...)
3. Trigger B16 is missing.

14:33 Online Split requests tape mount: UPH016 -->

15:30 Shift ends with smooth data taking in progress.

15:36 Online Split requests tape mount: UPM022 -->
KTEV1/tps3d2. (Heroic DAY shift does tape mount even
though it is after 15:30.

KTeV Logbook XXII, pages 168-176

Follow up Notes on Previous Shift

Items to Follow up on Next Shifts

1. NOTE (ECS): Please flip NM2SR polarity whenever NM4AN is changed from
POS to NEG (that's the next NM4AN flip).

2. JasonH requests that the shift crew page him IF beam is to be off for
more than an hour on a non-OWL shift.

Earl C. Swallow                           email:  earls@elmhurst.edu
Theophil W. Mueller Professor              (or)   earls@fnal.gov
   and Chairperson
Department of Physics                     Phone:  (630)-617-3577
Elmhurst College                          FAX:    (630)-617-3735
190 Prospect Avenue
Elmhurst, IL   60126-3296                 http://www.elmhurst.edu/~earls
"Books are lighthouses erected in the dark sea of time."
                                --  from Gargoyles  (The Scrolls of Merlin)